PS-IS-123/RC

 

MINUTES OF THE PS-ISOLDE COORDINATION MEETING

held on 1st October 2002

 

 

Present:

R. Catherall, S. de Man, A. Dorsival, V. Fedosseev, G.J. Focker, U. Georg, U. Koester, G. Le Scornet, J. Lettry, M. Lindroos, F. Locci, S. Marzari, S. Meunier, T. Nilsson, J. Parra-Lopez, J. Schipper, F. Wenander.

Excused:

D. Carminati, B. Crepieux

 

ISOLDE at URL: http://isolde.cern.ch

Consolidation projects at URL:

http://isolde-consolidation.web.cern.ch/isolde-consolidation/

PSMB meetings at URL:

http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html

 

1. Separators:

       The GPS was used for most of last week with measurements made on Ga and Cu. The next target change for a 6um Ta foil target will be on Wednesday 2nd October at 10:00.

       A UC2-C target #237 was placed on the HRS on Wednesday 25th September. Stable beam for the Tonnerre started on the following Sunday.

 

2. Beam lines:

       The users were not able to apply HT to LA0.BE10. Closer inspection revealed that there was a short circuit across a feed-through/connector inside the vacuum chamber. The short circuit may have occurred during a venting of the vacuum section while the HT remained. Due to the increasing amount of similar short circuits, S. de Man asks that the current of the power supplies be measured. J. Parra-Lopez will do this. When opening the beam line vacuum chambers, the following people should be contacted: J. Parra-Lopez (to ensure grounding of the HT), S. Meunier (for vacuum and o ring replacement) and A. Dorsival( radiation protection).

      

3. Tape Station:

On Friday 27th September, the tape at the tape station repeatedly dislodged itself from its reel. This problem is due to both the design of the tape station and the actual state of the tape. U. Koester also remarked that the displacement of the tape is slightly irregular and could lead to incorrect data. G.J. Focker will carry out repair work and replace the tape on Monday 7th October.

F. Locci will contact U. Koester to carry out tests on the tape station.

 

5. Vacuum:

       Both the NICOLE and MISTRAL beam line sections are now available to the vacuum group to carry out the modifications required when installing new turbo pumps. The vacuum group would like to remind everyone that it is no longer in a position to lend or give away material. All required vacuum material should be bought by the user and with the appropriate budget code.

 

6. AOB

       J. Schipper requests a converter target in order to repeat the conditions that provoked a delay in the 60kV recovery time earlier this year and to make further measurements. This may be possible in week 43…to be confirmed.

       U. Georg announced his departure from CERN at the end of this year. A technical engineer post has been opened for his replacement.

       Confirmation is given to S. Marzari to carry out the modifications on the target water circuit in an attempt to balance the water pressure/flow at each front end.

       The lack of cleaning in building 170 is becoming a problem. U. Georg will contact those responsible to ensure a regular cleaning.

       The users are asked to consider the needs of other groups when using material in the hall. When material is not being used it should be made available to others.

       Although the electrometers need replacing/repairing, it should be noted that the read-out can also be read from the console.

       M. Lindroos informs that the detailed planning of the shutdown work will be done next week. Those who would like to amend or inform of other pending tasks should do so as soon as possible.

 

Absences

T. Giles from 13.09 to 07.10

 

 

Next meeting:

Tuesday 8 October 2002 at 11.00 in the EP Conference Room A, building 304-1-001.